OpenStudy (anonymous):

whats the focus and directrix of x2=12y? or how do you find it?

terenzreignz (terenzreignz):

When it's already of the form \[\Large (x-h)^2 = b(y-k)\] Its focus would be the point \[\Large \left(h \ , \ k+\frac{b}{4 }\right)\] while its directrix would be the line \[\Large y =k-\frac{b}4 \]